# Identifying Dwarfs of MC Analog GalaxiEs (ID-MAGE): The Search for Satellites Around Low-Mass Hosts

### v3

@@ 1 Introduction
Dwarf galaxies are unique laboratories for studying the nature of dark matter and galaxy formation. Because of their shallow gravitational potentials, they are extremely sensitive to differences in cosmological and galaxy formation and evolution models, resulting in a range of testable predictions for galaxy formation. Dwarfs can also be dramatically changed by their environment. Currently, our understanding of dwarfs is predominantly based on observations of satellite galaxies around Milky Way (MW)-mass galaxies (e.g., Martin et al., 2013; Laevens et al., 2015; Koposov et al., 2015; Drlica-Wagner et al., 2015, 2020; Chiboucas et al., 2009, 2013; Crnojević et al., 2016, 2019; Müller et al., 2019; Smercina et al., 2018; Sand et al., 2014; Toloba et al., 2016; Mutlu-Pakdil et al., 2022, 2024; Bennet et al., 2019; Carlsten et al., 2022; Geha et al., 2017; Mao et al., 2024), which poses a risk of over-tailoring our models to observations of satellites systems of Milky Way-mass galaxies. Therefore, a statistical sample of satellites of dwarf galaxies is urgently needed.
@@
One particular point of interest is lower-mass host systems, which are testing grounds to study the effect of environment on dwarf galaxy formation (Wetzel et al., 2015), where gravitational effects are reduced. The standard Cold Dark Matter (CDM) model predicts that even moderate-sized dwarf galaxies should host their own small satellites (Munshi et al., 2019). The spatial clustering of 20 recently discovered ultra-faint (, ) galaxies provides evidence that the LMC itself might have fallen into the Local Group with its own satellite system (Patel et al., 2018; Battaglia et al., 2022). However, it also raises an interesting question about the luminosity function of the Magellanic association. The LMC has a fairly massive companion, the SMC, which is about 1.5 mag fainter than the LMC, but its next most luminous satellite is nearly 13 mag fainter (Hydrus 1, 4.7; Koposov et al. 2018). This results in a mag ‘gap’ in the satellite luminosity function. This large magnitude-gap is unexpected given the number of substructures expected for a DM halo containing a galaxy as luminous as the LMC. From simulations, satellites with are expected around an LMC-mass galaxy (e.g., Dooley et al. 2017a). M33 appears to be similarly lacking in bright satellites with just two known likely satellites with luminosities of (Martin et al., 2009; Collins et al., 2024). As M33 and the LMC are satellites of more massive hosts, their satellite systems may have been impacted by the MW/Andromeda.To fully understand the satellite population of LMC-mass galaxies and to test our models for populating hosts with satellites in general, it is essential to study the satellite systems of nearby LMC/SMC-mass galaxies. Although the Local Group is the only place to detect the extremely low luminosity satellites that are being found near the LMC, we can build up a sample of satellites with by targeting nearby low-mass galaxies.
@@
Recent efforts, including surveys like MADCASH (Magellanic Analog Dwarf Companions and Stellar Halos; Carlin et al. 2016), DELVE-DEEP (the DEEP component of DECam Local Volume Exploration Survey (DELVE); Drlica-Wagner et al. 2022), LBT-SONG (LBT Satellites of Nearby Galaxies Survey; Garling et al. 2021), and ELVES-Dwarf (Exploration of Local VolumE Satellites Dwarf survey; Li et al. 2025) have begun building up a sample of satellites around isolated dwarf galaxy hosts (Sand et al. 2015, 2024; Rich et al. 2012; Carlin et al. 2016, 2021, 2024; Davis et al. 2021, 2024; McNanna et al. 2024; Li et al. 2025). MADCASH, DELVE-DEEP and the nearby component of LBT-SONG focus on isolated hosts with stellar masses between . These programs focus on systems within Mpc, where dwarf satellites can be investigated using resolved stars from ground-based observations. Within this range, it is possible to push the discovery frontier of dwarf galaxies to fainter magnitudes, enabling a more profound understanding of their characteristics through their resolved stellar populations. ID-MAGE’s goal is to build a statistical sample of satellites of dwarf galaxies using integrated light searches to test dark matter and galaxy formation theories effectively. Surveys such as the Exploration of Local VolumE Satellites Survey (ELVES) and Satellites Around Galactic Analogs (SAGA) have demonstrated the value of building statistical samples of dwarf satellites around MW-mass hosts. ELVES utilized similar unresolved techniques to search for satellites of 30 MW-mass hosts out to a distance of 12 Mpc (Carlsten et al. 2022), while SAGA used spectroscopy to confirm satellites around 101 MW-mass hosts with distances between 25 and 41 Mpc (Geha et al. 2017; Mao et al. 2024).
@@
ID-MAGE aims to find unresolved satellites around 36 low-mass host galaxies with distances between 410 Mpc through a modified version of the well-established integrated light search algorithm (Bennet et al., 2017). ID-MAGE nearly triples the number of low-mass galaxies surveyed for satellites with well-characterized detection limits. Our satellite sample will provide vital clues for our understanding of galaxy evolution physics at these scales, and serve as a ground of comparison to the results obtained for the LMC and MW-mass hosts, yielding new insights into how host properties can affect satellite dwarf evolution. Additionally, with 36 low-mass hosts, we will reach sufficient statistical power to use their satellite populations as a test to the theoretical predictions (e.g., Dooley et al., 2017a; Nadler et al., 2022; Jahn et al., 2022) to constrain the physics that shapes the relationship between dwarf baryonic properties and DM halo mass.
@@
This paper presents the overview and initial results of the new survey ID-MAGE, including the candidate satellites identified, their photometric properties, and comparisons with cosmological predictions. In Section 2, we describe the goals of ID-MAGE and our host selections. Section 3 details the detection algorithm and the survey’s completeness. Section 4 details the photometry of the candidates. In Section 5, we compare our candidates to known Local Volume dwarf galaxies and compare the number of satellite candidates we detected to simulation predictions. Section 6 goes into our ongoing campaign to follow-up our candidate satellites. Finally, we summarize our key results in Section 7.
2 Survey Description
@@
A robust sample of satellite galaxies around hosts with a wide range of masses is required to test CDM models and our understanding of galaxy formation and evolution. The goal of ID-MAGE is to identify satellite galaxy systems around LMC/SMC-mass hosts in diverse environments. In this section, we present the selection of our hosts and their characteristics. We initially describe our host selection criteria in Section 2.1. Following that, we discuss the host environment in Section 2.2, where we describe the relative isolation of our hosts and highlight interesting environmental features. In Section 2.3, we summarize our use of the DESI Legacy Survey imaging data for ID-MAGE.
2.1 Host Selection
@@
We select our hosts from Cosmicflows-4 (Tully et al., 2023) –a catalog of the distances to 55,877 galaxies within z– choosing galaxies simply via a cut in luminosity ( Ks mag), distance ( Mpc), and Galactic latitude ( ). Our magnitude cuts translate to a host stellar mass range of 108.0 M⊙(one-fourth the SMC’s stellar mass of MM⊙) to M∗109.75 M⊙ (three times the LMC’s stellar mass of MM⊙), assuming M∗/LK =0.6 (McGaugh & Schombert 2014). We require TRGB distances for our hosts to accurately match the distances of the systems with confidence. We impose an absolute galactic latitude cut to avoid the plane of the MW. Additionally, we require coverage of the host galaxies out to a projected radius of 150 kpc in the DESI Legacy Imaging Surveys (Dey et al., 2019) Data Release (DR)-10 in the g- and r-bands. This ensures the full virial radius is covered by the imaging survey for each host. The hosts’ viral radii are estimated as 120140 kpc for LMC-mass galaxies and 80100 kpc for SMC-mass galaxies (Guo et al., 2013; Mutlu-Pakdil et al., 2021). Assuming Moster et al. (2010)’s stellar mass-to-halo mass relationship, the halo masses in log(Mhalo/M⊙) of our LMC-mass galaxies ranges from 11.15 to 11.40 and 10.70 to 11.05 for our SMC-mass hosts. We exclude galaxies that fall within the projected virial radius of a MW-mass galaxy and have either Mpc or kms-1 relative to the host. We also exclude host galaxies if there is a large known background galaxy cluster whose contamination would be difficult to remove from our satellite sample or if significant galactic cirrus impedes the search. In total, we exclude 35 potential hosts within the Legacy Survey for the reasons described above. This results in a total of 36 low-mass galaxies (9 in the LMC-mass range, 27 in the SMC-mass range). The complete list of hosts is presented in Table 1.
@@
Figure 1 demonstrates how our survey expands the mass range of host galaxies surveyed for satellite systems, with clearly defined completeness limits. Our target host galaxies are considerably less massive than those in ELVES, aligning more closely with the mass ranges surveyed in programs like LBT-SONG (Garling et al., 2021), MADCASH (Carlin et al., 2016, 2024), and DELVE-DEEP (Drlica-Wagner et al., 2022). Although most hosts are roughly LMC/SMC-like in stellar mass, our sample includes hosts both more and less massive than the LMC/SMC, which will allow insight into how host properties can affect satellite dwarf evolution. Unlike MADCASH and DELVE-DEEP, our selection criteria do not require the host galaxies to be isolated but they are required to not be satellites of larger galaxies (see Section 2.2).
@@ 2.2 Host Environment
Environment plays a significant role in the evolution of satellite galaxies. The weaker tidal and ram pressure interactions associated with lower-mass hosts may enable their satellites to preserve their neutral gas reservoirs and continue forming stars (e.g., Spekkens et al. 2014). To explore how environmental factors affect satellite evolution, our selected host galaxies vary in their proximity to other nearby galaxies, ranging from complete isolation with no other galaxies within 500 kpc, to being part of groups of dwarf galaxies.
@@
We utilize the catalogs from Cosmicflows-4 (Tully et al., 2009) and Kourkchi & Tully (2017) to check galaxies near our hosts. For each host, we select all galaxies within a projected radius of 450 kpc and with a distance in Cosmicflows-4 within 1 Mpc. Figure 2 showcases examples of three environments in our host sample. The surveyed areas are marked in gray. Galaxies with MKs are shown with their approximate virial radii indicated: 110 kpc for SMC-mass galaxies, 150 kpc for LMC-mass galaxies, and 300 kpc for more massive galaxies. As shown in Figure 2, our surveyed hosts are in diverse environments, ranging from those with no known galaxies with MKs within a projected radius of 450 kpc at a similar distance (left panel), to hosts that are part of larger galaxy groups (middle and right panels). See Appendix A for environment plots of other hosts.
@@
We also consider the environment of our hosts using the tidal index described in Karachentsev et al. (2013) (see Tables 1). The tidal index is a measure of the local stellar density derived from the stellar mass and distance of the nearest significant neighbors. accounts for the tidal contributions from the five most significant neighbors (Karachentsev et al., 2013). The majority of our hosts have low tidal index values, with 22 hosts having , indicating they are isolated galaxies. The other 15 have higher tidal indices, indicating that they are not isolated galaxies but instead occupy a range of field and group environments. The tidal indices mostly agree with environment plots in Appendix A, as hosts with no known neighbors have , while those that appear to be in denser environments have . A few notable exceptions are discussed in the appendix.
2.3 DESI Legacy Surveys Imaging Data
@@
We utilize the publicly available, wide-field DESI Legacy Imaging Surveys dr-10 data (Dey et al., 2019) for satellite candidate detection. For each host galaxy, we downloaded the -band data from the web server111https://www.legacysurvey.org/ covering a projected radius of 150 kpc in small parts (12′12′ regions) and reconstructed larger fields with the astropy (Astropy Collaboration et al., 2013, 2018) package reproject. We preferred the larger fields of view to run the dwarf detection algorithm more efficiently. The exact g-band depth varies across the survey area with an approximate 5 depth of g24.0. The data were retrieved with the native 0.262″ per pixel resolution and the ls-dr10 setting which merges the northern (MzLSBASS) and southern (DECam) imaging data at the declination of 32.375°.
@@
The coverage for each host extends to a projected radius of 150 kpc. This radius is slightly larger than the estimated 120140 kpc virial radius for LMC-mass galaxies, and it extends well beyond the estimated 80100 kpc virial radius of SMC-mass galaxies (Mutlu-Pakdil et al., 2021).
3 Dwarf Satellite Search
3.1 Algorithm Search
@@
We use a modified version of the dwarf detection algorithm described in Bennet et al. (2017) to identify candidate satellite galaxies. This algorithm is based on previous works (e.g., Dalcanton et al. 1997; Davies et al. 2016). The steps of the algorithm are illustrated in Figure 3 and briefly summarized below.
@@
The algorithm first masks foreground stars and known background galaxies in the Guide Star Catalog (GSC) 2.3.2 (Lasker et al., 2008) by creating a circular region covering the source and its outer halo. The mask region grows in size logarithmically based on the magnitude of the source. This initial step is similar to the masking procedure used in other low-surface brightness galaxy searches (e.g., van der Burg et al. 2016; Bennet et al. 2017; Carlsten et al. 2022). After bright source masking, Source Extractor (Bertin & Arnouts, 1996) is run on the masked image to identify sources with more than 25 pixels 5 above the sky level. These relatively bright sources, such as the centers of high surface brightness galaxies, are masked without attempting to remove their outer halos. The masked image is then spatially binned. Two versions of the binned image are created, one binned by 100100 pixels and the other by 5050 pixels. The binning corresponds to a spatial scale of 12001200 pc (600600 pc) at 10 Mpc and 500500 pc (250250 pc) at 4 Mpc. We choose these binning scales to detect large diffuse galaxies while remaining sensitive to smaller objects across the full distance range of the sample. We test different combinations of binning scales to optimize sensitivity and minimize false positives. As seen in Figure 3, our chosen binning scales enhances the detection of diffuse objects while effectively smoothing background variations. For both binning scales, pixels above 200 are masked to remove image artifacts such as chip gaps and diffraction spikes. Finally, we run the Source Extractor on the binned images, and pixels 4 above the background are cataloged. The two catalogs from the different binning scales are compared, and any objects found in both catalogs are forwarded for initial visual screening.
@@
To assess the survey completeness, we inject artificial galaxies into the data for each host galaxy using the Sersic2D function in astropy. These artificial galaxies are simulated with a Sérsic index of (Sersic, 1968), typical for diffuse dwarf galaxies (Koda et al., 2015; van Dokkum et al., 2015a; van der Burg et al., 2016). The simulated galaxies vary in absolute magnitudes from M7 to 13 (m1422.5 mag) and have effective radii (re) ranging from 2.6″ to 185″ (re=508720 pc). The artificial galaxies have a uniform distribution in apparent magnitude from m16.522.5 mag, with the number of artificial galaxies injected tapering off for m16.5. The artificial galaxies have a logarithmic distribution in effective radius with more smaller galaxies injected. The number of artificial dwarfs injected into each bin per host is available at Zenodo https://doi.org/10.5281/zenodo.15498887 (catalog doi:10.5281/zenodo.15498887) and Github222https://github.com/hunte22l/ID-MAGE_Completeness.git. For our detection efficiency tests, we model the artificial dwarfs as circular with zero ellipticity. As reported in van der Burg et al. (2016), moderate ellipticities do not impact the detection efficiency of the algorithm.
@@
As depicted in Figure 4, the algorithm’s detection efficiency is for larger (5.5″), brighter (20.2) and higher central surface brightness objects (26.026.5), with the efficiency rapidly falling to below 50% for galaxies with or for a Sérsic profile, depending on the host. The algorithm also loses sensitivity to very compact, bright objects, such as bright elliptical galaxies and stars, as they are masked out. The algorithm completeness drops off between 20.021.0. Based on Bennet et al. (2017) and Carlsten et al. (2022), we expect very few dwarf galaxies to lie within this region.
3.2 Visual Screening
This screening is conducted through a web interface that presents the cutouts of the original -band image, along with its masked, binned, and smoothed versions for each detected object. This display facilitates the easy identification of diffuse candidates and the removal of obvious false detections. To verify the effectiveness of our screening in identifying diffuse objects, we calibrate the display parameters using catalogs of known low-surface-brightness galaxies (SMUDGes; Systematically Measuring Ultra Diffuse Galaxies, Zaritsky et al. 2023), diffuse satellite galaxies (Bennet et al., 2017; Carlsten et al., 2022), and simulated galaxies injected into our science fields. These simulated galaxies are modeled using the astropy.modeling function Sersic2D as simple Sérsic profiles covering the full range of expected dwarf properties (m1622 mags, effective radius r 2.6″200″, Sérsic index 0.52.5, ellipticity of 00.4), and added before the masking process. We adjust the display settings for the screening to ensure consistent recovery of known low-surface-brightness galaxies and simulated diffuse galaxies. Simulated galaxies fail screening if they are obscured by galactic cirrus or superimposed on stars or other galaxies.

### v5

@@ I Introduction
Dwarf galaxies are unique laboratories for studying the nature of dark matter and galaxy formation. Because of their shallow gravitational potentials, they are extremely sensitive to differences in cosmological and galaxy formation and evolution models, resulting in a range of testable predictions for galaxy formation. Dwarfs can also be dramatically changed by their environment. Currently, our understanding of dwarfs is predominantly based on observations of satellite galaxies around Milky Way (MW)-mass galaxies (e.g., Martin2013; Laevens15; Koposov2015; Drlica2015; Drlica2020; Chiboucas09; Chiboucas2013; Crnojevic16; Crnojevic2019; Muller2019; Smercina18; Sand14; toloba16; BMP22; BMP24; Bennet19; ELVES; SAGAI; SAGA), which poses a risk of over-tailoring our models to observations of satellites systems of Milky Way-mass galaxies. Therefore, a statistical sample of satellites of dwarf galaxies is urgently needed.
@@
One particular point of interest is lower-mass host systems, which are testing grounds to study the effect of environment on dwarf galaxy formation (Wetzel2015), where gravitational effects are reduced. The standard Cold Dark Matter (CDM) model predicts that even moderate-sized dwarf galaxies should host their own small satellites (Munshi2019). The spatial clustering of 20 recently discovered ultra-faint (, ) galaxies provides evidence that the LMC itself might have fallen into the Local Group with its own satellite system (Patel18; Battaglia22). However, it also raises an interesting question about the luminosity function of the Magellanic association. The LMC has a fairly massive companion, the SMC, which is about 1.5 mag fainter than the LMC, but its next most luminous satellite is nearly 13 mag fainter (Hydrus 1, 4.7; Koposov18). This results in a mag ‘gap’ in the satellite luminosity function. This large magnitude-gap is unexpected given the number of substructures expected for a DM halo containing a galaxy as luminous as the LMC. From simulations, satellites with are expected around an LMC-mass galaxy (e.g., Dooley17). M33 appears to be similarly lacking in bright satellites with just two known likely satellites with luminosities of (Martin09; Collins24). As M33 and the LMC are satellites of more massive hosts, their satellite systems may have been impacted by the MW/Andromeda.To fully understand the satellite population of LMC-mass galaxies and to test our models for populating hosts with satellites in general, it is essential to study the satellite systems of nearby LMC/SMC-mass galaxies. Although the Local Group is the only place to detect the extremely low luminosity satellites that are being found near the LMC, we can build up a sample of satellites with by targeting nearby low-mass galaxies.
@@
Recent efforts, including surveys like MADCASH (Magellanic Analog Dwarf Companions and Stellar Halos; MADCASH), DELVE-DEEP (the DEEP component of DECam Local Volume Exploration Survey (DELVE); DELVE), LBT-SONG (LBT Satellites of Nearby Galaxies Survey; LBT-SONG), and ELVES-Dwarf (Exploration of Local VolumE Satellites Dwarf survey; ELVES-Dwarf) have begun building up a sample of satellites around isolated dwarf galaxy hosts (Sand15; sand24; Rich12; MADCASH; Carlin21; Carlin24; Davis2021; Davis2024; mcnanna24; ELVES-Dwarf). MADCASH, DELVE-DEEP and the nearby component of LBT-SONG focus on isolated hosts with stellar masses between . These programs focus on systems within Mpc, where dwarf satellites can be investigated using resolved stars from ground-based observations. Within this range, it is possible to push the discovery frontier of dwarf galaxies to fainter magnitudes, enabling a more profound understanding of their characteristics through their resolved stellar populations. ID-MAGE’s goal is to build a statistical sample of satellites of dwarf galaxies using integrated light searches to test dark matter and galaxy formation theories effectively. Surveys such as the Exploration of Local VolumE Satellites Survey (ELVES) and Satellites Around Galactic Analogs (SAGA) have demonstrated the value of building statistical samples of dwarf satellites around MW-mass hosts. ELVES utilized similar unresolved techniques to search for satellites of 30 MW-mass hosts out to a distance of 12 Mpc (ELVES), while SAGA used spectroscopy to confirm satellites around 101 MW-mass hosts with distances between 25 and 41 Mpc (SAGAI; SAGA).
@@
ID-MAGE aims to find unresolved satellites around 36 low-mass host galaxies with distances between 410 Mpc through a modified version of the well-established integrated light search algorithm (Bennet17). ID-MAGE nearly triples the number of low-mass galaxies surveyed for satellites with well-characterized detection limits. Our satellite sample will provide vital clues for our understanding of galaxy evolution physics at these scales, and serve as a ground of comparison to the results obtained for the LMC and MW-mass hosts, yielding new insights into how host properties can affect satellite dwarf evolution. Additionally, with 36 low-mass hosts, we will reach sufficient statistical power to use their satellite populations as a test to the theoretical predictions (e.g., Dooley17; Nadler2022; Jahn2022) to constrain the physics that shapes the relationship between dwarf baryonic properties and DM halo mass.
@@
This paper presents the overview and initial results of the new survey ID-MAGE, including the candidate satellites identified, their photometric properties, and comparisons with cosmological predictions. In Section II, we describe the goals of ID-MAGE and our host selections. Section III details the detection algorithm and the survey’s completeness. Section IV details the photometry of the candidates. In Section V, we compare our candidates to known Local Volume dwarf galaxies and compare the number of satellite candidates we detected to simulation predictions. Section VI goes into our ongoing campaign to follow-up our candidate satellites. Finally, we summarize our key results in Section VII.
II Survey Description
@@
A robust sample of satellite galaxies around hosts with a wide range of masses is required to test CDM models and our understanding of galaxy formation and evolution. The goal of ID-MAGE is to identify satellite galaxy systems around LMC/SMC-mass hosts in diverse environments. In this section, we present the selection of our hosts and their characteristics. We initially describe our host selection criteria in Section II.1. Following that, we discuss the host environment in Section II.2, where we describe the relative isolation of our hosts and highlight interesting environmental features. In Section II.3, we summarize our use of the DESI Legacy Survey imaging data for ID-MAGE.
II.1 Host Selection
@@
We select our hosts from Cosmicflows-4 (Tully23) –a catalog of the distances to 55,877 galaxies within z– choosing galaxies simply via a cut in luminosity ( Ks mag), distance ( Mpc), and Galactic latitude ( ). Our magnitude cuts translate to a host stellar mass range of 108.0 M⊙(one-fourth the SMC’s stellar mass of MM⊙) to M∗109.75 M⊙ (three times the LMC’s stellar mass of MM⊙), assuming M∗/LK =0.6 (McGaugh14). We require TRGB distances for our hosts to accurately match the distances of the systems with confidence. We impose an absolute galactic latitude cut to avoid the plane of the MW. Additionally, we require coverage of the host galaxies out to a projected radius of 150 kpc in the DESI Legacy Imaging Surveys (Dey19) Data Release (DR)-10 in the g- and r-bands. This ensures the full virial radius is covered by the imaging survey for each host. The hosts’ viral radii are estimated as 120140 kpc for LMC-mass galaxies and 80100 kpc for SMC-mass galaxies (Guo13; BMP2021). Assuming Moster10’s stellar mass-to-halo mass relationship, the halo masses in log(Mhalo/M⊙) of our LMC-mass galaxies ranges from 11.15 to 11.40 and 10.70 to 11.05 for our SMC-mass hosts. We exclude galaxies that fall within the projected virial radius of a MW-mass galaxy and have either Mpc or kms-1 relative to the host. We also exclude host galaxies if there is a large known background galaxy cluster whose contamination would be difficult to remove from our satellite sample or if significant galactic cirrus impedes the search. In total, we exclude 35 potential hosts within the Legacy Survey for the reasons described above. This results in a total of 36 low-mass galaxies (9 in the LMC-mass range, 27 in the SMC-mass range). The complete list of hosts is presented in Table 1.
@@
Figure 1 demonstrates how our survey expands the mass range of host galaxies surveyed for satellite systems, with clearly defined completeness limits. Our target host galaxies are considerably less massive than those in ELVES, aligning more closely with the mass ranges surveyed in programs like LBT-SONG (LBT-SONG), MADCASH (MADCASH; Carlin24), and DELVE-DEEP (DELVE). Although most hosts are roughly LMC/SMC-like in stellar mass, our sample includes hosts both more and less massive than the LMC/SMC, which will allow insight into how host properties can affect satellite dwarf evolution. Unlike MADCASH and DELVE-DEEP, our selection criteria do not require the host galaxies to be isolated but they are required to not be satellites of larger galaxies (see Section II.2).
@@ II.2 Host Environment
Environment plays a significant role in the evolution of satellite galaxies. The weaker tidal and ram pressure interactions associated with lower-mass hosts may enable their satellites to preserve their neutral gas reservoirs and continue forming stars (e.g., spekkens14). To explore how environmental factors affect satellite evolution, our selected host galaxies vary in their proximity to other nearby galaxies, ranging from complete isolation with no other galaxies within 500 kpc, to being part of groups of dwarf galaxies.
@@
We utilize the catalogs from Cosmicflows-4 (EDD) and Kourkchi17 to check galaxies near our hosts. For each host, we select all galaxies within a projected radius of 450 kpc and with a distance in Cosmicflows-4 within 1 Mpc. Figure 2 showcases examples of three environments in our host sample. The surveyed areas are marked in gray. Galaxies with MKs are shown with their approximate virial radii indicated: 110 kpc for SMC-mass galaxies, 150 kpc for LMC-mass galaxies, and 300 kpc for more massive galaxies. As shown in Figure 2, our surveyed hosts are in diverse environments, ranging from those with no known galaxies with MKs within a projected radius of 450 kpc at a similar distance (left panel), to hosts that are part of larger galaxy groups (middle and right panels). See Appendix A for environment plots of other hosts.
@@
We also consider the environment of our hosts using the tidal index described in Karachen13 (see Tables 1). The tidal index is a measure of the local stellar density derived from the stellar mass and distance of the nearest significant neighbors. accounts for the tidal contributions from the five most significant neighbors (Karachen13). The majority of our hosts have low tidal index values, with 22 hosts having , indicating they are isolated galaxies. The other 15 have higher tidal indices, indicating that they are not isolated galaxies but instead occupy a range of field and group environments. The tidal indices mostly agree with environment plots in Appendix A, as hosts with no known neighbors have , while those that appear to be in denser environments have . A few notable exceptions are discussed in the appendix.
II.3 DESI Legacy Surveys Imaging Data
@@
We utilize the publicly available, wide-field DESI Legacy Imaging Surveys dr-10 data (Dey19) for satellite candidate detection. For each host galaxy, we downloaded the -band data from the web server111https://www.legacysurvey.org/ covering a projected radius of 150 kpc in small parts (12′12′ regions) and reconstructed larger fields with the astropy (astropy13; astropy18) package reproject. We preferred the larger fields of view to run the dwarf detection algorithm more efficiently. The exact g-band depth varies across the survey area with an approximate 5 depth of g24.0. The data were retrieved with the native 0.262″ per pixel resolution and the ls-dr10 setting which merges the northern (MzLSBASS) and southern (DECam) imaging data at the declination of 32.375°.
@@
The coverage for each host extends to a projected radius of 150 kpc. This radius is slightly larger than the estimated 120140 kpc virial radius for LMC-mass galaxies, and it extends well beyond the estimated 80100 kpc virial radius of SMC-mass galaxies (BMP2021).
III Dwarf Satellite Search
III.1 Algorithm Search
@@
We use a modified version of the dwarf detection algorithm described in Bennet17 to identify candidate satellite galaxies. This algorithm is based on previous works (e.g., Dalcanton97; Davies16). The steps of the algorithm are illustrated in Figure 3 and briefly summarized below.
@@
The algorithm first masks foreground stars and known background galaxies in the Guide Star Catalog (GSC) 2.3.2 (GSC) by creating a circular region covering the source and its outer halo. The mask region grows in size logarithmically based on the magnitude of the source. This initial step is similar to the masking procedure used in other low-surface brightness galaxy searches (e.g., vanderburg16; Bennet17; ELVES). After bright source masking, Source Extractor (Sextractor) is run on the masked image to identify sources with more than 25 pixels 5 above the sky level. These relatively bright sources, such as the centers of high surface brightness galaxies, are masked without attempting to remove their outer halos. The masked image is then spatially binned. Two versions of the binned image are created, one binned by 100100 pixels and the other by 5050 pixels. The binning corresponds to a spatial scale of 12001200 pc (600600 pc) at 10 Mpc and 500500 pc (250250 pc) at 4 Mpc. We choose these binning scales to detect large diffuse galaxies while remaining sensitive to smaller objects across the full distance range of the sample. We test different combinations of binning scales to optimize sensitivity and minimize false positives. As seen in Figure 3, our chosen binning scales enhances the detection of diffuse objects while effectively smoothing background variations. For both binning scales, pixels above 200 are masked to remove image artifacts such as chip gaps and diffraction spikes. Finally, we run the Source Extractor on the binned images, and pixels 4 above the background are cataloged. The two catalogs from the different binning scales are compared, and any objects found in both catalogs are forwarded for initial visual screening.
@@
To assess the survey completeness, we inject artificial galaxies into the data for each host galaxy using the Sersic2D function in astropy. These artificial galaxies are simulated with a Sérsic index of (sersic68), typical for diffuse dwarf galaxies (koda15; vandokkum15; vanderburg16). The simulated galaxies vary in absolute magnitudes from M7 to 13 (m1422.5 mag) and have effective radii (re) ranging from 2.6″ to 185″ (re=508720 pc). The artificial galaxies have a uniform distribution in apparent magnitude from m16.522.5 mag, with the number of artificial galaxies injected tapering off for m16.5. The artificial galaxies have a logarithmic distribution in effective radius with more smaller galaxies injected. The number of artificial dwarfs injected into each bin per host is available at Zenodo https://doi.org/10.5281/zenodo.15498887 (catalog doi:10.5281/zenodo.15498887) and Github222https://github.com/hunte22l/ID-MAGE_Completeness.git. For our detection efficiency tests, we model the artificial dwarfs as circular with zero ellipticity. As reported in vanderburg16, moderate ellipticities do not impact the detection efficiency of the algorithm.
@@
As depicted in Figure 4, the algorithm’s detection efficiency is for larger (5.5″), brighter (20.2) and higher central surface brightness objects (26.026.5), with the efficiency rapidly falling to below 50% for galaxies with or for a Sérsic profile, depending on the host. The algorithm also loses sensitivity to very compact, bright objects, such as bright elliptical galaxies and stars, as they are masked out. The algorithm completeness drops off between 20.021.0. Based on Bennet17 and ELVES, we expect very few dwarf galaxies to lie within this region.
III.2 Visual Screening
This screening is conducted through a web interface that presents the cutouts of the original -band image, along with its masked, binned, and smoothed versions for each detected object. This display facilitates the easy identification of diffuse candidates and the removal of obvious false detections. To verify the effectiveness of our screening in identifying diffuse objects, we calibrate the display parameters using catalogs of known low-surface-brightness galaxies (SMUDGes; Systematically Measuring Ultra Diffuse Galaxies, SMUDGESV), diffuse satellite galaxies (Bennet17; ELVES), and simulated galaxies injected into our science fields. These simulated galaxies are modeled using the astropy.modeling function Sersic2D as simple Sérsic profiles covering the full range of expected dwarf properties (m1622 mags, effective radius r 2.6″200″, Sérsic index 0.52.5, ellipticity of 00.4), and added before the masking process. We adjust the display settings for the screening to ensure consistent recovery of known low-surface-brightness galaxies and simulated diffuse galaxies. Simulated galaxies fail screening if they are obscured by galactic cirrus or superimposed on stars or other galaxies.